Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by hakyology » Sun Jul 06, 2008 11:50 am

There is a totla of 120 marbles in a box, each of which is red, green, blue or white. If one of marble is drawn from the box at random, the probability that it will be white is 1/4 and the probability that it will be green is 1/3. What is the probability that the marble will be either red or blue?

A) 1/6
B) 1/4
C) 2/7
D) 1/3
E) 5/12

Please explain. Thanks.


The answer is E.
Join the discussion
Source: — Problem Solving |

by sibbineni » Sun Jul 06, 2008 11:54 am
Given Data:
Total no of marbles:120
The probability of white is 1/4==>30 White
The probability of green is 1/3==>40 Green
so the rest will be either red or blue

120-30-40==>50 red or blue

so the probability of red or blue is 50/120===>5/12
Join the discussion

by andes1 » Sun Jul 06, 2008 1:59 pm
The easy way
RED GREEN WHITE BLUE
?------1/3--------1/4------?

OK. NOW MAKE A SUM
1/3 + 1/4 …. = 7/12
This is the combined probability of green and white.

Now only make a sustracction from the 100% or total that is 12/12, or simply 1

1- 7/12 = 5/12
That is the probability of blue and red combined.
LEARNING ENGLIS H
Join the discussion